# PhysGaia: A Physics-Aware Dataset of Multi-Body Interactions for Dynamic Novel View Synthesis

### v1

@@ Abstract
We introduce PhysGaia, a novel physics-aware dataset specifically designed for Dynamic Novel View Synthesis (DyNVS), encompassing both structured objects and unstructured physical phenomena. Unlike existing datasets that primarily focus on photorealistic reconstruction, PhysGaia is created to actively support physics-aware dynamic scene modeling. Our dataset provides complex dynamic scenarios with rich interactions among multiple objects, where they realistically collide with each other and exchange forces. Furthermore, it contains a diverse range of physical materials, such as liquid, gas, viscoelastic substance, and textile, which moves beyond the rigid bodies prevalent in existing datasets. All scenes in PhysGaia are faithfully generated to strictly adhere to physical laws, leveraging carefully selected material-specific physics solvers. To enable quantitative evaluation of physical modeling, our dataset provides essential ground-truth information, including 3D particle trajectories and physics parameters, e.g., viscosity. To facilitate research adoption, we also provide essential integration pipelines for using state-of-the-art DyNVS models with our dataset and report their results. By addressing the critical lack of datasets for physics-aware modeling, PhysGaia will significantly advance research in dynamic view synthesis, physics-based scene understanding, and deep learning models integrated with physical simulation—ultimately enabling more faithful reconstruction and interpretation of complex dynamic scenes. Our datasets and codes are available in the project website [^1] http://cvlab.snu.ac.kr/research/PhysGaia.
@@ 1 Introduction
(a) Liquid with FLIP solver: hanok (left) and ship (right)
(b) Gas with Pyro solver: box-smoke (left) and pisa (right)
(c) Viscoelastic substances with MPM solver: pancake (left) and jelly party (right)
(d) Textile with Vellum solver: lucy (left) and basin (right)
Figure 1: Examples from the proposed physics-aware dataset, PhysGaia. They exhibit complex physical interactions between multiple objects composed of diverse materials such as liquid, gas, viscoelastic substance, and textile. This dataset will foster physics reasoning in dynamic scenes.
@@
Since the introduction of Neural Radiance Fields (NeRF) [1], the Novel View Synthesis (NVS) algorithms based on deep learning have progressed rapidly. While early research primarily targeted static scenes, recent efforts have extended the scope to dynamic scene reconstruction to better support diverse and interactive AR/VR applications. This emerging direction, known as Dynamic Novel View Synthesis (DyNVS), aims to reconstruct 4D scenes from input videos and synthesize photorealistic images at novel viewpoints and time steps that are not seen during training.
@@
The evolution of DyNVS has been closely tied to the development of suitable datasets. Early DyNVS datasets [2, 3, 4] primarily provided multiview training videos and typically involved only limited object motion. These datasets laid the foundation for early DyNVS research [3, 5, 6, 7, 8, 9, 10, 11], which focused on modeling scene dynamics by deforming canonical geometries over time. Recently, to better support AR/VR applications, datasets captured using handheld mobile devices have been introduced [7, 12, 13]. These datasets consist of monocular videos, and recent research has primarily focused on mitigating overfitting under such limited supervision. Building on these recent datasets, subsequent works [14, 15, 16, 17, 18, 19, 20] also have commonly extended recent Gaussian Splatting [21] to dynamic scenarios, aiming to achieve photorealistic rendering. While these efforts have led to significant improvements in visual quality, photorealism alone is insufficient for presenting the underlying dynamics and physical properties of real-world scenes.
@@
As DyNVS continues to advance, a natural next step is to move beyond photorealism and incorporate physical realism—enabling models not only to render how scenes look, but also to reason about how they behave. Recent pioneering works [22, 23, 24, 25, 26] have begun to explore this direction by integrating differentiable physics simulation into Gaussian Splatting [21] frameworks. Despite this growing interest, current research largely remains limited to 4D generation [27, 28, 29, 30] or simplified DyNVS scenarios [26, 31], often restricted to single object or single material. Consequently, more complex scenarios—such as multi-object interactions or the modeling of diverse physical materials—are still significantly under-explored. In response, we introduce PhysGaia, a new benchmark designed to support and accelerate research in these emerging and challenging directions.
Our contributions are summarized as follows:
@@
  • We introduce PhysGaia, a physics-aware dataset featuring rich interactions among multiple objects and encompassing a wide range of physical materials, including liquids, gases, textiles, and viscoelastic substances, as illustrated in Figure 1.
@@
  • We evaluate existing state-of-the-art DyNVS methods on PhysGaia, revealing their fundamental limitations in achieving physical realism and demonstrating the potential for improvement in this field.
The rest of this paper is organized as follows. Section 2 reviews related work and highlights the uniqueness of our dataset. We discuss the main properties and potential research enabled by our dataset in Section 3. Section 4 details material-specific physics solvers used for dataset construction and methodologies for generating multi-body interaction scenarios. Section 5 shows our analysis, including evaluations of recent DyNVS algorithms on PhysGaia, We conclude the paper in Section 6.
@@ 2.1 Dynamic Novel View Synthesis
In recent years, significant advances have been made in novel view synthesis [1, 32, 33, 34, 35, 21]. Although this field initially have focused on static scene reconstruction but now it extended to handling dynamic scenes, which is known as Dynamic Novel View Synthesis (DyNVS). Early DyNVS methods were built on Neural Radiance Fields (NeRF) [5, 6, 7, 8, 9, 10, 11], which usually modeled scene dynamics either by implictly modeling with temporal inputs [5, 6] or directly estimating the time-wise deformation of canonical geometry through auxiliary neural networks [7, 8, 9, 10, 11]. Following the emergence of 3D Gaussian Splatting (3DGS) [21], recent DyNVS research [17, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47] has shifted toward Gaussian Splatting-based representations to achieve higher fidelity and real-time inference. Since Gaussian Splatting is an explicit representation, additional deformation networks can directly modify the scale, rotation, and position of each gaussian primitive over time.
In a significant development, research is now moving towards constraining the deformation of gaussian primitives to adhere to physical realism. PhysGaussian [22] pioneered this by combining an MPM (Material Point Method) [48] simulator with Gaussian Splatting [21], treating each gaussian primitive as a particle within the MPM simulation. This work has inspired many subsequent studies [31, 26, 27, 28, 29, 30] integrating physics-aware priors into Gaussian Splatting; however, these efforts remain largely confined to generation tasks [27, 28, 29, 30], with only a few addressing Dynamic Novel View Synthesis (DyNVS) [31, 26]. Even for DyNVS, methods are typically limited to single objects or single materials, with most focusing on viscoelastic substances. Consequently, exploring physics-aware DyNVS research involving rich object interactions and diverse physical materials remains significantly underexplored. We believe our dataset can serve as a foundational resource for this crucial direction.
@@
Table 1: Comparison with Dynamic Novel View Synthesis datasets. This table summarizes key attributes including the number of total scenes having test videos for DyNVS evaluation, the number of scenes with physical phenomena, presence of multi-body interaction, ground-truth physics parameters, e.g., viscosity, and ground-truth 3D particle trajectories. Our PhysGaia dataset uniquely offers a comprehensive set of diverse scenes with complex multi-body interactions and complete ground-truth physical information, a significant advancement over prior benchmarks.
Datasets Scene Stats Physics Information Capture Type
# total scene # physics Interaction Physics param. 3D Traj.
Plenoptic [2]   6       1(gas) No No No Multiview
D-NeRF [3]   8       1(viscoelastic) No No No Monocular
NVIDIA Dynamic [4]   8       0 No No No Monocular
Nerfies [7]   4       0 No No No Monocular
HyperNeRF [12]   4       0 No No No Monocular
DyCheck [13]   7       1(textile) No No No Monocular
NeRF-DS [49]   8       0 No No No Monocular
EvDNeRF [50]   6       0 No No No Multiview
Synthetic Soccer [51]   3       0 No No No Multiview
HDR-HexPlane [52]   8       0 No No No Multiview
[…]
Table 1 provides a comparative overview of our dataset alongside existing multiview and monocular DyNVS datasets. The initial DyNVS datasets [2, 3, 4] primarily employed multiview configurations or captured scenes with very limited motion, typically involving mostly rigid objects. For example, Plenoptic [2] was captured with sufficiently many multiview cameras at fixed positions. To advance DyNVS toward practical AR/VR applications, more user-friendly datasets, often captured using handheld mobile phones, were later introduced. Nerfies [7] pioneered handheld iPhone captures, though its scenes remained largely static. HyperNeRF [12] then introduced more rapid and varied motions, and DyCheck [13] further resolved camera teleportation issues seen in HyperNeRF. These monocular datasets have motivated research mitigating overfitting to training videos in these settings. Some approaches [14, 15, 16] focus on leveraging additional priors such as diffusion models [53], depth estimation models [54], and point trackers [55], while others [15, 17, 18, 19, 20] emphasize constraining deformation by adopting techniques like motion factorization [56] or As-Rigid-As-Possible (ARAP) regularization [57]. Nonetheless, the primary objective across these datasets remains photorealistic reconstruction, with limited emphasis on physics-aware dynamic modeling  [^2] Note that some recent datasets have also been introduced, but they are typically tailored to specific scenarios, such as event cameras [50], HDR rendering [52], specular lighting effects [49], or human motion capture [51]..
Unlike existing DyNVS datasets, our PhysGaia dataset offers scenes with multi-body interactions involving diverse physical materials, as detailed in Table 1. While some datasets [2, 3, 13] include limited scenes with physical phenomena, these are typically few in number, not generated using accurate physics solvers, and lack complex multi-object interactions. In this context, PhysGaia occupies a unique position and holds strong potential to spur advancements in physics-aware dynamic scene modeling and reasoning.
@@
Table 2: Comparison with physics-simulated datasets. “Multi-body Inter.” denotes multi-body interaction. Unlike existing datasets, our PhysGaia offers rich interactions, a diverse range of physical materials, and access to accurate physics parameters and ground-truth 3D particle trajectories, making it a uniquely valuable resource for advancing physics-based understanding of dynamic scenes.
Datasets Multi-body Inter. Materials Physics Information
Liquids Gas Textile Viscoelastic Physics params. 3D traj.
@@
ScalarFlow [64] No $\checkmark$ Yes Yes
[…]
(a) PAC-NeRF [63]
(b) Spring-Gaus [31]
(c) ScalarFlow [64]
Figure 2: Visualization of datasets most similar to our PhysGaia. While all of these datasets address physical phenomena, they are limited in several key aspects: limited coverage of physical materials, overly simplified dynamics, and an absence of rich multi-body interactions.
@@
To provide a comprehensive comparison, we evaluate our PhysGaia against other 4D datasets generated via physics simulation, as summarized in Table 2. Since DyNVS tasks require multiview RGB imagery, we focus on datasets that offer such data. While some existing datasets support multi-body interactions, they often lack rich interaction dynamics or are confined to a single material type. For instance, clothed human datasets [58, 59, 60] exhibit textile-human interactions but are limited to textiles and human-centric motions. In terms of ground-truth physical information, a few datasets [63, 31, 64] provide both physical parameters and 3D trajectories. However, these datasets are typically restricted to single materials and do not capture complex, multi-body interactions, as illustrated in Figure 2. Specifically, PAC-NeRF [63] (Figure 2(a)) focuses on liquids and viscoelastic materials, but its liquid scenarios are highly viscous, behaving similarly to viscoelastic materials. Spring-Gaus [31] (Figure 2(b)) is restricted to viscoelasticity, and ScalarFlow [64] (Figure 2(c)) is limited to non-interacting gases.
In stark contrast, our PhysGaia directly addresses these limitations. It provides rich interactions, a diverse range of physical materials, and access to accurate simulation parameters and ground-truth trajectories. This makes PhysGaia a uniquely valuable dataset for advancing our understanding of physics in dynamic scenes.
3 Dataset Properties and Research Impact
@@
We propose PhysGaia to advance physically realistic reconstruction in Dynamic Novel View Synthesis (DyNVS), extending the conventional reconstruction task beyond photorealism. As highlighted in Tables 1 and 2 of Section 2, our dataset uniquely features complex multi-body interactions involving a diverse range of physical materials, distinguishing it within the DyNVS landscape. We believe this comprehensive dataset holds significant potential to enhance the understanding of physics in dynamic scenes. This section outlines the key properties of our dataset and discusses potential future research directions it enables.
3.1 Dataset Properties
Complex physics-aware dynamics
(a) Multi-body interaction: Jelly party
(b) Multi-body interaction: Ship
(c) Splashing (non-locally rigid)
(d) Specular modeling
(e) Refraction
Figure 3: Visualization of physics properties in PhysGaia. Alongside multi-body interactions, PhysGaia also includes various physical phenomena like splashing, refraction, and specular effects.
Our PhysGaia dataset comprises 17 scenes that include multi-body interactions. Figure 1 illustrates a subset of these scenes. We further visualize the detailed examples in Figure 3(a) and Figure 3(b), which exhibit collisions between viscoelastic objects causing rapid changes in motion direction, and rigid-liquid surface impacts resulting in splashing, respectively. To ensure that these scenes adhere to physical laws with accurately calculated force exchange among multi bodies, we carefully select material-specific solvers: FLIP for liquids, Pyro for gases, MPM for viscoelastic materials, and Vellum for textiles. Further details on simulation configurations for handling multi-body interactions are found in Section 4.
Beyond multi-body interactions, our dataset also exhibits various physical phenomena such as non-locally rigid motion (common in liquid and gas scenes), specular reflection, and refraction, as shown in Figures 3(c)3(d), and 3(e), respectively. These properties enhance the realism of our dataset and enable a wide range of downstream tasks and research applications.
Ground-truth physics information

### v2

# PhysGaia: A Physics-Aware Benchmark with Multi-Body Interactions for Dynamic Novel View Synthesis

@@ Abstract
We introduce PhysGaia, a novel physics-aware benchmark for Dynamic Novel View Synthesis (DyNVS) that encompasses both structured objects and unstructured physical phenomena. While existing datasets primarily focus on photorealistic appearance, PhysGaia is specifically designed to support physics-consistent dynamic reconstruction. Our benchmark features complex scenarios with rich multi-body interactions, where objects realistically collide and exchange forces. Furthermore, it incorporates a diverse range of materials, including liquid, gas, textile, and rheological substance, moving beyond the rigid-body assumptions prevalent in prior work. To ensure physical fidelity, all scenes in PhysGaia are generated using material-specific physics solvers that strictly adhere to fundamental physical laws. We provide comprehensive ground-truth information, including 3D particle trajectories and physical parameters (e.g., viscosity), enabling the quantitative evaluation of physical modeling. To facilitate research adoption, we also provide integration pipelines for recent 4D Gaussian Splatting models along with our dataset and their results. By addressing the critical shortage of physics-aware benchmarks, PhysGaia can significantly advance research in dynamic view synthesis, physics-based scene understanding, and the integration of deep learning with physical simulation, ultimately enabling more faithful reconstruction and interpretation of complex dynamic scenes.
Figure 1: Visualization of key properties of PhysGaia. Unlike existing benchmarks limited to single objects or materials, PhysGaia is a physics-aware benchmark featuring complex multi-body interactions across diverse substances (liquids, gases, rheological materials, and textiles). By providing ground-truth 3D trajectories and physical parameters, it uniquely enables the evaluation of physical realism alongside traditional photorealism. In addition to multi-body collisions, our dataset captures (c) splashing effects characterized by non-local rigid motion, as well as (d–e) complex optical phenomena such as specular reflection and refraction.
footnotetext: indicates equal contribution.
@@ 1 Introduction
Since the emergence of Neural Radiance Fields (NeRF), Novel View Synthesis (NVS) algorithms based on deep learning have witnessed rapid advancements. While initial research primarily focused on static environments, recent efforts have shifted toward dynamic scene understanding to meet the demands of immersive and interactive AR/VR applications. This emerging direction, termed Dynamic Novel View Synthesis (DyNVS), aims to reconstruct complex 4D spatiotemporal scenes from video inputs, enabling the synthesis of photorealistic images at novel viewpoints and time steps beyond the training distribution.
@@
The evolution of DyNVS has been closely related to the availability of diverse benchmarks. Early DyNVS datasets [40, 57, 80] primarily provided multi-view training videos, albeit with relatively constrained object motion. These benchmarks established the groundwork for initial research [57, 17, 12, 55, 8, 16, 61, 15], which focused on modeling temporal dynamics by deforming a canonical geometry over time. More recently, benchmarks featuring videos captured with handheld mobile devices have been introduced [55, 56, 18]. These datasets consist of monocular sequences, prompting research to focus on mitigating overfitting under such sparse supervision. Building upon these foundations, subsequent works [33, 70, 84, 28, 36, 35, 7, 32] have extended Gaussian Splatting (GS) [31] to dynamic scenarios, aiming for high-fidelity, photorealistic rendering.
@@
As DyNVS continues to advance, a natural next step is to move beyond photorealism and incorporate physical realism, enabling models not only to render how scenes look, but also to reason about how they behave. Recent pioneering works [75, 83, 48, 4, 29] have begun exploring this direction by integrating differentiable physics simulation into the 4D Gaussian Splatting (4DGS) framework. Despite this burgeoning interest, existing research remains largely confined to 4D generation or simplified DyNVS scenarios [29, 85], often restricted to isolated objects with a single material type. Consequently, more complex environments—characterized by multi-body interactions and a diverse range of physical materials such as liquids and gases—remain significantly underexplored. Furthermore, while real-world videos of such complex scenarios can be readily captured, they inherently lack the ground-truth physical information (e.g., 3D trajectories and material parameters) necessary for a rigorous and quantitative evaluation of physical realism. In response, we introduce PhysGaia, a new benchmark with novel evaluation metrics designed to support and accelerate research in these challenging directions. Figure 1 illustrates examples with complex physical interactions in our benchmark. Our contributions are summarized as follows:
@@
  • We introduce PhysGaia, a physics-aware dataset featuring rich interactions among multiple objects and encompassing a wide range of physical materials, including liquids, gases, textiles, and rheological substances.
@@
  • We test existing DyNVS methods on PhysGaia, revealing their fundamental limitations in achieving physical realism and demonstrating the potential for improvement in this field.
@@ 2.1 Dynamic Novel View Synthesis
In recent years, significant advances have been made in novel view synthesis [52, 34, 9, 20, 69, 54, 31]. Although this field initially focused on static scene reconstruction, it has since extended to dynamic scenes, known as Dynamic Novel View Synthesis (DyNVS). Early DyNVS methods were built on Neural Radiance Fields (NeRF) [17, 12, 55, 8, 16, 61, 15], which usually modeled scene dynamics either by implicitly modeling with temporal inputs [17, 12] or directly estimating the time-wise deformation of canonical geometry through auxiliary neural networks [55, 8, 16, 61, 15]. Following the emergence of 3D Gaussian Splatting (3DGS)[31], recent DyNVS research has shifted toward Gaussian-based representations, leading to the development of 4D Gaussian Splatting (4DGS) [28, 79, 74, 42, 45, 49, 22, 43, 38, 13, 68, 47, 65, 32]. In 4DGS, an additional deformation network is employed to animate canonical Gaussian primitives over time, enabling efficient and high-quality modeling of dynamic scenes.
Modeling dynamic scenes with 4DGS is increasingly moving toward incorporating physical laws to govern motion. PhysGaussian [75] pioneers this direction by combining an MPM [66] simulator with Gaussian Splatting, treating each Gaussian primitive as a particle within the particle-grid simulation framework of MPM. This work has inspired many subsequent studies [85, 29, 30, 46, 27, 58] integrating physics-aware priors into 4DGS; however, these efforts remain largely confined to generation tasks [30, 46, 27, 58], with only a few addressing Dynamic Novel View Synthesis (DyNVS) [85, 29]. Even among DyNVS methods, existing approaches are typically limited to single-object scenes with homogeneous materials, predominantly focusing on rheological substances. Consequently, physics-aware DyNVS involving rich object interactions and diverse physical materials remains largely underexplored, a gap our dataset is designed to address.
@@
Table 1: Comparison with Dynamic NVS datasets. Our PhysGaia benchmark offers diverse scenes with complex multi-body interactions and complete ground-truth physical information (parameters and trajectories).
Datasets Scenes Physics Info. View
Total Phys. Inter. Param. Traj.
Plenoptic [40] 6 1(G) Multi
NVIDIA Dynamic [80] 8 0 Mono
FluidNexus [19] 240 240(G) Multi
DyCheck [18] 7 1(T) Mono
Nerfies [55] 4 0 Mono
HyperNeRF [56] 4 0 Mono
NeRF-DS [76] 8 0 Mono
EvDNeRF [2] 6 0 Multi
Synthetic Soccer [39] 3 0 Multi
HDR-HexPlane [73] 8 0 Multi
D-NeRF [57] 8 1(R) Mono
Phystwin [29] 22 22(R) Multi
PhysGaia (Ours) 17 17 Both
G: Gas, T: Textile, R: Rheological
Table 1 provides a comparative overview of our dataset alongside existing DyNVS datasets. The early DyNVS datasets [40, 57, 80] primarily employed multiview configurations or captured scenes with very limited motion, typically involving mostly rigid objects. These datasets paved the way for DyNVS research, with subsequent works [57, 17, 12, 55, 8, 16, 61, 15] exploring deforming canonical geometries over time to model scene dynamics. To advance DyNVS toward practical AR/VR applications, more user-friendly datasets captured using handheld mobile phones were later introduced. Nerfies [55] pioneered handheld iPhone captures, though its scenes remained largely static. HyperNeRF [56] then introduced more rapid and varied motions, and DyCheck [18] further addressed camera teleportation artifacts observed in HyperNeRF. These monocular datasets have motivated research aimed at mitigating overfitting to training views, with some approaches [33, 84, 70] leveraging additional priors such as diffusion models [60], depth estimators [77], and point trackers [78], while others [28, 70, 35, 36, 7] constrain deformation via motion factorization [67] or As-Rigid-As-Possible (ARAP) regularization [64]. Nonetheless, the primary objective across these datasets remains photorealistic reconstruction, with limited emphasis on physics-aware dynamic modeling.
Unlike existing DyNVS datasets, PhysGaia provides scenes featuring multi-body interactions across diverse physical materials, as shown in Table 1. Although some datasets [40, 57, 18, 19, 29] include scenes with physical phenomena, they are limited in scale, lack accurate physics simulation, or do not capture complex multi-object interactions. In this regard, PhysGaia occupies a unique position with strong potential to advance physics-aware dynamic scene modeling.
@@
Table 2: Comparison with physics-simulated datasets for the DyNVS task. Our PhysGaia benchmark offers diverse materials, multi-body interactions, and complete physical ground truth (parameters and trajectories).
Datasets Multi-body Materials Physics Info.
Interactions L G T R Params. 3D Traj.
@@
PhysGaia (Ours) Yes $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Yes Yes
L: Liquid, G: Gas, T: Textile, R: Rheological
To provide a comprehensive comparison, we also compare PhysGaia with existing 4D datasets generated via physics simulation, as summarized in Table 2. Since DyNVS tasks require multiview RGB imagery, we focus on datasets that provide such data.
@@
Some existing datasets support multi-object interactions, but either lack rich interaction dynamics or are limited to a single material type. For instance, clothed human datasets [81, 86, 71] naturally exhibit textile-body interactions, yet are restricted to textile materials and human-centric motions. In terms of ground-truth physical information, a few datasets [41, 85, 14] provide both physical parameters and 3D trajectories; however, they remain confined to single-material settings and do not capture rich multi-object interactions. PAC-NeRF [41] focuses on liquids and viscoelastic materials, but its liquid scenarios are constrained to highly viscous flows, making them behaviorally similar to viscoelastic materials. Spring-Gaus [85] is restricted to viscoelasticity, while ScalarFlow [14] concentrates on gas. In contrast, PhysGaia directly addresses these limitations by offering rich multi-object interactions, a diverse range of physical materials, and accurate simulation parameters and trajectories, making it a uniquely valuable resource for advancing physics-aware understanding of dynamic scenes.
(a) Liquid with FLIP solver: hanok (left) and ship (right)
(b) Gas with Pyro solver: box-smoke (left) and pisa (right)
(c) Rheological substances with MPM solver: pancake (left) and jelly party (right)
(d) Textile with Vellum solver: lucy (left) and basin (right)
Figure 2: Examples from the proposed physics-aware benchmark, PhysGaia. They exhibit complex physical interactions between multiple objects composed of diverse materials such as liquid, gas, rheological substance, and textile. More importantly, our benchmark enables evaluation of physics realism to foster physics reasoning in dynamic scenes.
3 PhysGaia
@@
We propose PhysGaia to advance physically realistic reconstruction in DyNVS, moving beyond mere photorealism. As shown in Tables 1 and 2, PhysGaia uniquely features complex multi-body interactions across a diverse range of physical materials, distinguishing it from existing DyNVS datasets. We believe PhysGaia holds significant potential to deepen the understanding of physics in dynamic scenes.
3.1 Benchmark Properties
Complex scenarios with physics-aware dynamics
Our dataset consists of 17 scenes featuring multi-object interactions, as visualized in Figure 2. To ensure that each scene adheres to physical laws with accurately calculated force exchange among objects, we carefully select material-specific solvers: Fluid-Implicit Particle (FLIP) [5] for liquids, Pyro [62] for gases, Vellum [63] for textiles, and Material Point Method (MPM) [66] for rheological materials. Beyond multi-body interactions, our dataset also captures various physical phenomena such as non-locally rigid motion commonly observed in liquid and gas scenes, specular reflection, and refraction, as shown in Figure 1. These properties enhance the realism of our dataset and support a wide range of downstream tasks and research applications.
Providing physics parameters